OpenStudy (anonymous):

Solve the following system of equations. −0.7x+0.6y=2.3 0.5x+1.8y=−6.1 help will medal

OpenStudy (anonymous):

x= y=

OpenStudy (anonymous):

use a graphics calculator and graph them? And find its point of intersection.

OpenStudy (anonymous):

or you can do it via algebraic means. depends what is required

OpenStudy (anonymous):

ty your awesome @chris00

OpenStudy (anonymous):

: )

OpenStudy (anonymous):

what was your answer?

OpenStudy (anonymous):

to graph it on your calc, we tend to get y on its own.

OpenStudy (anonymous):

\[y _{1}=\frac{ 2.3+0.7x }{ 0.6 }\] \[y _{2}=\frac{ -6.1-0.5x }{ 1.8 }\]

OpenStudy (anonymous):

just graph these equation on your calculator and find its point on intersection

OpenStudy (anonymous):

you should get (-5,-2)

OpenStudy (anonymous):

You could also use a matrix and apply cramer's rule. and as @chris00 said you should get X = -5 , Y = -2

OpenStudy (anonymous):

depends if @Missiey is required to do it. I usually go the most quickest method if the question doesn't tell me to use a specific method

OpenStudy (anonymous):

thanks chtis

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

@chris00 ty i aced the test

OpenStudy (anonymous):

legend